A survey was conducted to assess the market share of Ringtone, a telecommunications company. Out of a random sample of 200 people who have mobile phone contracts surveyed, 38 of them have Ringtone contracts.

(a) Calculate the sample proportion of people who have Ringtone contracts.

(b) Ringtone has established a target market share of at least 15%. Use the sample proportion found in (a), test whether this target has been attained at a 5% significance level.

(c) Use the information found in (a) to set up a 99% confident interval for the population proportion of Ringtone’s market share. Make clear all assumptions that are needed to justify this calculation.

(d) Suppose that before the survey was undertaken, Ringtone wanted to ensure that the 90% confident interval (CI) for the proportion of people who have Ringtone contracts that they end up with a margin of error of no more than 5%. How many people should it have sought to survey?
Answer #1

a) \widehat p = 38/200 = 0.19

b) H0: P > 0.15

    H1: P < 0.15

The test statistic z = (\widehat p- P)/sqr(P(1 - P)/n)

                              = (0.19 - 0.15)/sqrt(0.15 * (1 - 0.15)/200)

                              = 1.58

P-value = P(Z < 1.58)

             = 0.9429

Since the P-value is greater than the significance level (0.9429 > 0.05), so we should not reject the null hypothesis.

So at 5% significance level there is sufficient evidence to conclude that the target of at least 15% has been attained.

c) At 99% confidence interval the critical value is z0.005 = 2.58

The 99% confidence interval for population proportion is

\widehat p +/- z0.005 * sqrt(\widehat p(1 - \widehat p)/n)

= 0.19 +/- 2.58 * sqrt(0.19 * (1 - 0.19)/200)

= 0.19 +/- 0.072

= 0.118, 0.262

d) At 90% confidence interval the critical value is z0.05 = 1.645

Margin of error = 0.05

or, z0.05 * sqrt(P(1 - P)/n) = 0.05

or, 1.645 * sqrt(0.5 * 0.5/n) = 0.05

or, n = (1.645 * sqrt(0.5 * 0.5)/0.05)^2

or, n = 271